# VBART: The Turkish LLM

### 02 Mar 2024 v1

@@ Abstract
We present VBART, the first Turkish sequence-to-sequence Large Language Models (LLMs) pre-trained on a large corpus from scratch. VBART are compact LLMs based on good ideas leveraged from BART and mBART models and come in two sizes, Large and XLarge. Fine-tuned VBART models surpass the prior state-of-the-art results in abstractive text summarization, title generation, text paraphrasing, question answering and question generation tasks. They allow fine-tuning for future text generation tasks and datasets, carving a new path for Turkish Natural Language Processing (NLP) research. Our work shows that having a pre-trained LLM for Turkish outperforms up to 3x multilingual models, improving existing results and providing efficient models for training and inference. Moreover, we show that our monolingual tokenizer is 7x more efficient than OpenAI’s multilingual tokenizer. Last but not least, we introduce a method to enlarge an existing pre-trained LLM and question the relevancy of Chinchilla Scaling Law to sequence-to-sequence masked language models. Our fine-tuned models, tokenizer and cleaned web corpus of 135 GB are publicly available at huggingface.co/vngrs-ai.
@@ 1.   Introduction
Thus, we present VBART, the first dedicated seq2seq LLM pre-trained for Turkish on a large scale. First, we train VBART-Large from scratch on a cleaned corpus of 135.7 GBs of Turkish text based on the pre-training task of BART and the model architecture of mBART. Then, we create VBART-XLarge by enlarging the prior model by doubling the number of encoder and decoder layers, using the weights of VBART-Large. Both models surpass the previous state-of-the-art in abstractive text summarization, paraphrasing, question answering and question generation tasks. Our contributions to Turkish NLP research are as follows:
@@
    135 GBs of cleaned text
@@ 3.1.   Tokenizer
Before training a model from scratch, a tokenizer is needed first. We train SentencePiece Kudo and Richardson (2018) Unigram Model Tokenizer with a vocabulary size of 32,000 on random subsets of OSCAR Abadji et al. (2022), OPUS Zhang et al. (2020) and Wikipedia dump corpora. Training on 10 GBs of text required 500+ GBs of memory and took 3 hours on 96 CPU cores.
@@ 3.4.   Training Corpus
Training corpus is made of Turkish sections of OSCAR-2201 Abadji et al. (2022) and mC4 Xue et al. (2020), which contain 10.8M and 87.7 million pages, respectively. They are concatenated to obtain the final training corpus of 98.5M pages in total. However, web-crawled data is often noisy and full of keywords, titles and other non-sentence texts. In order to obtain a higher-quality dataset, these pages are cleaned using a chain of rules and heuristics. See the Appendix for the details of the data cleaning process. The cleaned corpus holds 135.7 GB of space on disk, contains 50.3M pages and is made of 25.33B subword tokens in total.
@@ 3.8.   Implementation
SentencePiece Kudo and Richardson (2018) library is used to train the Unigram Model tokenizer. The network, data generator and training are implemented using Tensorflow Abadi et al. (2015) framework. Moreover, trained networks and the tokenizer are converted to corresponding Huggingface implementations in order to use Huggingface Transformers’ Wolf et al. (2019)text generation utilities. They can be accessed on the Huggingface hub  [^1] https://huggingface.co/vngrs-ai along with the pre-training dataset.
@@ 4.3.   Text Paraphrasing
VBART models are fine-tuned for 20 and 25 epochs, respectively, on a mixture of OpenSubtitles Lison and Tiedemann (2016), TED Cettolo et al. (2012) and Tatoeba Artetxe and Schwenk (2019) datasets mentioned in the paper published by Alkurdi et al. (2022). The original paper suggests a method to improve the dataset quality. Since the filtered data is not published, our models are trained and evaluated on the unfiltered data. Then, we compare our models to the models that were fine-tuned on unfiltered data. Results are evaluated using BERTScore, BLEU, Rouge-L, Meteor and Translation Error Rate (TER) metrics. BERTScore is computed using cased 222dbmdz/bert-base-turkish-cased last reached in 19 Oct 2023 and uncased 333dbmdz/bert-base-turkish-uncased last reached in 19 Oct 2023 versions. Results are reported in Table 3. Except for TER, which is the lower, the better, it is the higher, the better for all metrics.
@@ 4.4.   Question Generation and Answering
Then we evaluate on the test splits of TQuADv1, TQuADv2 and XQuAD Artetxe et al. (2019) and report the results on Table4. BLEU-1, BLEU2 and Rouge-L are computed for question generation task. F1 and Exact Match (EM) are computed for question answering task. In order to be comparable, we do not drop the intersected test samples from the training dataset, following the original work. Therefore, these results are useful for comparison only and are not realistic enough for the tasks themselves.
@@ 5.1.   Tokenizer
Tokenizers of multilingual models are trained on a number of languages, and consequently, their vocabulary size is often large. For instance, OpenAI’s BPE tokenizer, which is used with ChatGPT and GPT-4 models, has a vocabulary size of 100,277. This results in a non-optimal tokenizer for a low-resource language such as Turkish. As a result, it takes 3.10 tokens to encode a Turkish word on average for the OpenAI tokenizer, while it takes 1.33 for our tokenizer, whose vocabulary size is one-third of the prior. Consequently, our tokenizer is $\frac{3.10}{1.33}\times\frac{100K}{32K}=7.28$ times more compact in terms of representation power.
@@ 6.   Conclusion
In this work, we presented the first sequence-to-sequence LLMs for the Turkish language. All of the components, dataset, tokenizer and models are prepared and trained for Turkish from scratch. We showed our tokenizer’s ability of compact representation over multilingual ones, compared the two models proposed in this work, introduced a method to enlarge an existing LLM and questioned the relevancy of Chinchilla Scaling Law to encoder-decoder models. Achieving new state-of-the-art results for abstractive text summarization, paraphrasing, question answering and question generation, we showed that monolingual pre-trained models surpass multilingual ones when fine-tuned for downstream tasks, despite having significantly fewer parameters.

### 14 Mar 2024 v2

@@ Abstract
We present VBART, the first Turkish sequence-to-sequence Large Language Models (LLMs) pre-trained on a large corpus from scratch. VBART are compact LLMs based on good ideas leveraged from BART and mBART models and come in two sizes, Large and XLarge. Fine-tuned VBART models surpass the prior state-of-the-art results in abstractive text summarization, title generation, text paraphrasing, question answering and question generation tasks. They allow fine-tuning for future text generation tasks and datasets, carving a new path for Turkish Natural Language Processing (NLP) research. Our work shows that having a pre-trained LLM for Turkish outperforms up to 3x multilingual models, improving existing results and providing efficient models for training and inference. Moreover, we show that our monolingual tokenizer is up to 11x more efficient than multilingual tokenizers. Last but not least, we introduce a method to enlarge an existing pre-trained LLM and question the relevancy of Chinchilla Scaling Law to sequence-to-sequence masked language models. Our fine-tuned models, tokenizer and cleaned vngrs-web-corpus of 135 GB are publicly available at huggingface.co/vngrs-ai.
@@ 1.   Introduction
Thus, we present VBART, the first dedicated seq2seq LLM pre-trained for Turkish on a large scale. First, we train VBART-Large from scratch on a cleaned corpus of 135.7 GBs of Turkish text based on the pre-training task of BART and the model architecture of mBART. Then, we create VBART-XLarge by enlarging the prior model by doubling the number of encoder and decoder layers, using the weights of VBART-Large. Both models surpass the previous state-of-the-art in abstractive text summarization, paraphrasing, title generation, question answering and question generation tasks. Our contributions to Turkish NLP research are as follows:
@@
    vngrs-web-corpus that consist of 135 GBs of cleaned text
@@ 3.1.   Tokenizer
Before training a model from scratch, a tokenizer is needed first. We use SentencePiece Kudo and Richardson (2018) Unigram Model Tokenizer that is part of VNLP Turker et al. (2024) library. It was trained on 10 GB of text that consists of random subsets of OSCAR Abadji et al. (2022), OPUS Zhang et al. (2020) and Wikipedia dump corpora. Tokenizer training required 500+ GBs of memory and took 3 hours on 96 CPU cores.
@@ 3.4.   Training Corpus
Training corpus is made of Turkish sections of OSCAR-2201 Abadji et al. (2022) and mC4 Xue et al. (2020), which contain 10.8M and 87.7 million pages, respectively. They are concatenated to obtain the final training corpus of 98.5M pages in total. However, web-crawled data is often noisy and full of keywords, titles and other non-sentence texts. In order to obtain a higher-quality dataset, these pages are cleaned using a chain of rules and heuristics. See the Appendix for the details of the data cleaning process. The cleaned corpus holds 135.7 GB of space on disk, contains 50.3M pages and is made of 25.33B subword tokens in total. We call this curated dataset vngrs-web-corpus.
@@ 3.8.   Implementation
SentencePiece Kudo and Richardson (2018) library is used to train the Unigram Model tokenizer. The network, data generator and training are implemented using Tensorflow Abadi et al. (2015) framework. Moreover, trained networks and the tokenizer are converted to corresponding Huggingface implementations in order to use Huggingface Transformers’ Wolf et al. (2019)text generation utilities. They can be accessed on the Huggingface hub  [^1] https://huggingface.co/vngrs-ai along with the vngrs-web-corpus.
@@ 4.2.   Title Generation
Figure 2: Relative number of tokens spent to encode Turkish text in average. The numbers below show the vocabulary size of each tokenizer. Despite their large vocabulary size, multilingual tokenizers spend up to 96% more tokens than VBART Tokenizer.
@@ 4.3.   Text Paraphrasing
VBART models are fine-tuned for 20 and 25 epochs, respectively, on a mixture of OpenSubtitles Lison and Tiedemann (2016), TED Cettolo et al. (2012) and Tatoeba Artetxe and Schwenk (2019) datasets mentioned in the paper published by Alkurdi et al. (2022). The original paper suggests a method to improve the dataset quality. Since the filtered data is not published, our models are trained and evaluated on the unfiltered data. Then, we compare our models to the models that were fine-tuned on unfiltered data. Results are evaluated using BERTScore, BLEU, Rouge-L, Meteor and Translation Error Rate (TER) metrics. BERTScore is computed using cased 222dbmdz/bert-base-turkish-cased last reached in 19 Oct 2023 and uncased 333dbmdz/bert-base-turkish-uncased last reached in 19 Oct 2023 versions. Results are reported in Table 3. Except for TER, which is the lower, the better; for all metrics, the higher, the better.
@@ 4.4.   Question Generation and Answering
Then we evaluate on the test splits of TQuADv1, TQuADv2 and XQuAD Artetxe et al. (2019) and report the results on Table4. BLEU-1, BLEU2 and Rouge-L are computed for question generation task. F1 and Exact Match (EM) are computed for question answering task. In order to be comparable, we do not drop the intersected test samples from the training dataset, following the original work. Also, note that following Çağatay Akyön et al. (2022), we report metrics up to a single decimal in order to be comparable. Therefore, these results are useful for comparison only and are not realistic enough for the tasks themselves.
@@ 5.1.   Tokenizer
Although one often thinks of the number of model parameters when comparing computational efficiency, it is equally crucial to be able to represent text using fewer tokens. Tokenizers of multilingual models are trained on multilingual corpora. This has two consequences. First, their vocabulary size is chosen to be large to be comprehensive enough. Second, despite their large vocabulary size, it takes more tokens to encode a text in a language compared to a tokenizer that was trained in the specific language alone.
We compare VBART Tokenizer to mBART, mT5 and OpenAI’s cl100k_base tokenizers. In doing so, we randomly sample 1M pages from vngrs-web-corpus and tokenize each page. Then we count the total number of tokens generated by each tokenizer. Figure 2 shows the relative number of tokens required to encode Turkish text, normalized by VBART Tokenizer. Despite having almost 8x vocabulary size, mBART and mT5 tokenizers spend 19.70% and 40.79% more tokens, respectively. OpenAI’s cl100k_base tokenizer, which is used along with Chat-GPT and GPT-4 models, spends 95.99% more tokens despite its more than 3x vocabulary size.
Moreover, taking vocabulary sizes into account, we can claim that VBART Tokenizer’s representation power is $\frac{119.70}{100}\times\frac{250,054}{32,000}=9.35$ times mBART’s. This is 11.00 and 6.14 for mT5 and OpenAI tokenizers, respectively.
@@ 6.   Conclusion
In this work, we presented the first sequence-to-sequence LLMs for the Turkish language. All of the components; dataset, tokenizer and models are prepared and trained for Turkish from scratch. We showed our tokenizer’s ability of compact representation over multilingual ones, compared the two models proposed in this work, introduced a method to enlarge an existing LLM and questioned the relevancy of Chinchilla Scaling Law to encoder-decoder models. Achieving new state-of-the-art results for abstractive text summarization, paraphrasing, title generation, question answering and question generation, we showed that monolingual pre-trained models surpass multilingual ones when fine-tuned for downstream tasks, despite having significantly fewer parameters.
@@ 7.   Bibliographical References
  • Turker et al. (2024) Meliksah Turker, Erdi Ari, and Aydin Han. 2024. Vnlp: Turkish nlp package. arXiv preprint arXiv:2403.01309.